A. ethan needs to determine the number and type of roots for the equation using the discriminant. his solution is shown. explain what ethan did correctly and what he did incorrectly and then correct his mistakes
1 b^2-4ac=-5^2-4(1)(10)=-25-40=-65
2 since -65<0 there are 2 non real solution
B. ethan decides to complete the square to rewrite the quadratic in vertex form and find the solutions to the equation when y=0. his solution is shown.identify what ethan did incorrectly and then explain what he should have done instead.(note: identify each error once)
1. y-10=x^2-5x+10-10-->y-10=x^2-5x
2. -5^2=-25
3.y-10-25=x^2-5x-25
4. y-35=(x-5)^2
5. y=(x-5)^2-35
6. 0=(x-5)^2-35
7. (x-5)^2=35
8.x-5=35/2 or x-5=-35/2
9. x=35/2+10/2 or x=-35/2+10/2
10. x=45/2 or x=-25/2

c. explain whether ethane solutions match his answer to part A

Answer by richwmiller(17219)

b^2-4ac=-5^2-4(1)(10)=-25-40=-65
(-5)^2=+25
should be +25-40=-15
interpretation stays
b)again
(-5)^2=+25
c)his answers are real we need complex
x = 1/2 (5+i sqrt(15))
x = 1/2 (5-i sqrt(15))